Label the points O(0, 0), A(-2, 4), and B(1, 1) on the graph
of f (x) x 2. Determine the images of these points on the transformed function. Label the images O',A' and B'

f(x)=-(x-2)^2
1
Expert's answer
2020-12-06T10:28:14-0500

Label the points O(0, 0), A(-2, 4), and B(1, 1) on the graph

of f (x) =x2


transformation function

f(x) = -(x-2)2


f(0) = -(0-2)2 = -4

O' (0,-4)


f(-2) = -(-2-2)2 = -16

A' (-2,-16)


f(1) = -(1-2)2 = -1

B' (1,-1)